OpenStudy (anonymous):

how do u do this.. i for got how plz help! 4x - 10 ------------ -2

OpenStudy (anonymous):

- 5/4 is the answer I hope I helped.

OpenStudy (anonymous):

how did u get the answer...

OpenStudy (e.mccormick):

You need to get x on one side of the equation.

OpenStudy (anonymous):

isnt it already?

OpenStudy (e.mccormick):

Well, if you do not have it equal to something, it is a bit hard to know what the solution should be. If there is nothing other than the fraction....

OpenStudy (e.mccormick):

\[\frac{ 4x-10 }{-2 }\] is all you have?

OpenStudy (anonymous):

Graphical form: simplifies to Text form: 4*x-10/-2 simplifies to 4*x+5 Look at 4x-10/(-2). Removed extra sign in front of It becomes .4*x+5 Result: 4*x+5

OpenStudy (anonymous):

I tried

OpenStudy (anonymous):

:l ...can u like teach me how to do this step by step? i compleately for got how to do this

OpenStudy (anonymous):

@e.mccormick ya thats the whole question..

OpenStudy (e.mccormick):

If it says to graph it, solve for y, or something else, then there are steps. Same for simplify, and so on. But with just an equation and no directions of what to do with it....

OpenStudy (anonymous):

it says Use the distributive property to simplify the expression below:

OpenStudy (e.mccormick):

Aaaah. Well, write the fraction as teo seperate fractions, then simplify.

OpenStudy (anonymous):

so.. 4x ova -2 & -10 ova -2??

OpenStudy (e.mccormick):

Yes. Now, simlify those.

OpenStudy (e.mccormick):

\[\frac{ 4x }{-2 }-\frac{ 10 }{-2 }\] is basically where you are at. So what is next...

OpenStudy (anonymous):

this is gunna sound stupid but how do u simplify that?

OpenStudy (e.mccormick):

Well, you have fractions with a 2 in the bottom. So what is half of th top numbers? Then do not forget the negative sign.

OpenStudy (anonymous):

-5 ova -1?

OpenStudy (anonymous):

for the right..

OpenStudy (e.mccormick):

Which also means? Two negatives makes a?

OpenStudy (anonymous):

+?

OpenStudy (e.mccormick):

Yes.

OpenStudy (e.mccormick):

If you think about that part like division, a -1 gopes into -5 all of five times, so the negative goes away. How is the left part going?

OpenStudy (anonymous):

2x ova -1??

OpenStudy (e.mccormick):

Yes. So what happens to the negative there?

OpenStudy (anonymous):

i dnt kno

OpenStudy (e.mccormick):

Let me put it another way. What is anuthing over 1? So \[\frac{ 6 }{1 }\] or \[\frac{ N }{1 }\] or anything.

OpenStudy (anonymous):

2 over 1? now uv confuzzed me >_<

OpenStudy (e.mccormick):

Sry bout that. It is like dividing by 1. Anything divided by 1 is itself. BUT, you have a negative, and you can not just get rid of the negative, so it has to stay. That help?

OpenStudy (anonymous):

no...but so the answr is 

OpenStudy (e.mccormick):

That is not quite the simplified form. Very, very close. You just need to deal with the -1s.

OpenStudy (anonymous):

how

OpenStudy (e.mccormick):

Above we talked about how \[\frac{ -5 }{ -1 } = 5\]

OpenStudy (anonymous):

so 2x = 5??

OpenStudy (e.mccormick):

So if we do at leas that much, we get \[\frac{ 2x }{ -1 }+5\] Now what do you think happens to the left part?

OpenStudy (anonymous):

were do u get the + from

OpenStudy (e.mccormick):

Not quite. You lost the negative. In multiplication or division, the nuber f negatives is key to if you have a negative or a positive. Two negatives is a positive, Three negatives is a negatrive. So on and so forth. 54 negatives is positive because there is an even number. So if I were dividing or mutiplying all of 54 numbers and they were all negatives....

OpenStudy (anonymous):

ok i get it now...uhh thx?

OpenStudy (e.mccormick):

Kk. I was going to say http://www.mathsisfun.com/multiplying-negatives.html might explain it better. So, what do you have for a final answer when you keep the one negative and change the other to a positive?